OpenStudy (tacomaco90):

PLEASE HELP ME ASAP Bricks are delivered to a work site and stacked in rows and columns, forming a rectangular prism. The length of the prism is 1 foot greater than its width, and its height is 2 feet less than its width. Find the dimensions of the prism formed by the bricks, given that its volume is 40 cubic feet.

OpenStudy (ashash146):

h=w-2, l=w+1, l*w*h=40 (w+1)*w*(w-2)=40 h = 2, l = 5, w = 4